# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v1

@@ Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains using test data. A highly effective CTTA method involves applying layer-wise adaptive learning rates, and selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. In this work, we aim to overcome these limitations by identifying layers through the quantification of model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity in order to approximate the domain shift, followed by adjusting their learning rates accordingly. Overall, this approach leads to a more robust and stable optimization than prior approaches. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C and demonstrate the efficacy of our method against standard benchmarks and prior methods.
@@ 1 Introduction
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [6, 44]. Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [44]. While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation: since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s pre-trained knowledge due to continuous adaptations and large parameter updates across tasks.
@@
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In [32, 8, 44, 37], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA [45] includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA [41] introduces the attachment of "meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
@@
More recently, the authors in [21] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be "surgically" or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) [33] was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) [36] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
@@
In this paper, we present a novel method called PALM to push the limit of adaptive learning rate methods for CTTA. Our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions. Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the KL divergence between the model output and a uniform distribution. This approach eliminates dependence on pseudo-labels and enables a more accurate estimation of layer-wise importance. Subsequently, the learning rates of the parameters of these layers are adjusted based on the degree of domain shift observed in the test data. We utilize parameter sensitivity as an indicator of the degree of distribution shift and employ weighted moving averages to aggregate sensitivity from previous data batches. Compared to existing adaptive learning rate methods, our proposed approach offers a better estimation of domain shift for adjusting learning rates.
To summarise, our contributions are as follows:
  • We present PALM, a novel approach to continual test-time adaptation. Specifically, we demonstrate improved schemes for adaptive learning rate methods for CTTA.
  • Considering the prediction uncertainty of the model at test-time, we design an approach to automatically select layers that would require further adaptation and push the learning rates of parameters of these layers for model updates, by estimating the domain shift based on parameter sensitivity.
  • Our method delivers promising results in both continual and gradual TTA across three standard benchmark datasets. In addition, we conduct and demonstrate ablation experiments on the different components of our work.
@@ 2 Related Works
Test-time adaptation (TTA): The overarching target of TTA is to adapt source models to unlabelled test data in real-time in a source-free online training fashion [16, 42, 44]. TENT [44] optimizes the Batch Normalisation (BN) parameters by minimizing the Shannon entropy [40] of the pseudo labels. In [37], the statistics of the BN parameters are replaced by the ones of the input corrupted data. Adacontrast [3] combines contrastive learning and utilizes the memory module to update the entire model’s parameters. While previous TTA methods showed significant improvements, they were demonstrated for a single domain shift at a time [4, 26, 37, 44] and did not consider continual distribution shifts.
Continual TTA (CTTA): CTTA poses an adaptation challenge wherein the model must continually adjust to evolving target data, accommodating a sequence of dynamic domain shifts. The first work, CoTTA [45], considers addressing this challenging task by setting up a teacher-student network, with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with forgetting and accumulation of errors with time. EcoTTA [41] introduces "meta networks" i.e. a BN layer and a convolution block are attached to the frozen source model. At every stage, an $L_{1}$ regularisation loss between the outputs of the meta network and that of the source model is computed to handle catastrophic forgetting.
Efficient fine-tuning approaches: Freezing model parameters to learn information from downstream data has witnessed a good body of work [39, 27, 51, 22, 7, 9]. Depending upon the distribution shift of the target task, in [21], the authors demonstrated that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive-learning rate methods: Regarding the studies on the importance of each layer in training, [47, 49] concluded that deeper layers extract task-specific features, as opposed to the shallow layers. Based on this finding, AutoLR [35] design an adapted learning rate (LR) scheme based on the weight variations between layers. Despite attempts to set different initial learning rates for each layer, these rates remain mostly constant throughout training. In the context of CTTA, LAW [33] used the FIM [36] to calculate the domain shift and automatically adjust the learning rates of parameters. However, to obtain the layer importance, the pseudo-labels are used to compute the FIM, which would be inaccurate. In addition, a direct accumulation of FIMs across domains would lead to error accumulation. Furthermore, the LR "weight" of each layer is upper-bounded by the base LR that is used, which ignores the relative "sensitivity" of parameters towards adaptation of the current batch and domain.
Parameter sensitivity: In the literature about neural network pruning, understanding the sensitivity of parameters towards training has been studied well [19, 28, 43, 29, 23]. The overall idea is that parameters with lower sensitivity need to be adapted more with an increased learning rate, and vice-versa. [23], in addition, introduces a temporal variation of sensitivity, due to uncertainty, as another factor during the training of models for natural language understanding and machine translation tasks.
3 Background
@@ 3.1 Problem Setting
The objective of continual test-time adaptation (CTTA) is to adapt a fixed discriminative pre-trained source model, parameterized by ${\theta_{s}}$ and trained on the source dataset ($\mathcal{X}^{source}$, $\mathcal{Y}^{source}$), to multiple incoming tasks $\mathcal{T}_{i}$ = {$x_{k}$}${}_{k=1}^{n}$ where $\mathcal{T}_{i}$ refers to the $i^{th}$ task and n is the total number of batches of $\mathcal{T}_{i}$, with changes in domain distributions between tasks. At each time step, ${\theta_{s}}$ is gradually adapted to each incoming batch ($x_{k}$) of task $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta}$. In this entire process of adaptation, the source dataset is unavailable for use, due to privacy and storage constraints. As evident, the source model is first parameterized by ${\theta_{s}}$ before the adaptation begins. With the continuum of tasks, at time step t, the complete model parameters ${\theta}^{t}$ consists of a matrix of parameters, given by ${\theta}^{t}$ = ${\theta^{t}_{1}}$, $\theta^{t}_{2}$,…, ${\theta^{t}_{n}}$,…., ${\theta^{t}_{N}}$}, where ${\theta^{t}_{n}}$, at time t, denotes the parameters of the $n^{th}$ trainable layer out of a total of N.
3.2 Analysis of Adaptive Learning Rate Schemes for CTTA
@@
In this subsection, we provide a detailed view and analysis of surgical fine-tuning [21], AutoLR [35], and LAW [33]. This is essential to understand the design principles and reasons for having an adaptive learning rate scheme for CTTA, which also includes the motivation of our work.
@@
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution, surgical fine-tuning [21] empirically show that heuristically fine-tuning only the first convolutional block results in a better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
Inspired by [47, 49], AutoLR [35] claims that the weight changes across layers should be in increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$, at epoch e, is defined as,
$\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$ (1)
@@
They hypothesize that weight changes in the initial layers should be small because they extract general features whereas the deeper layers extract task-specific features. A constrained adaptive LR optimization scheme is set up to force the weight variations to increase with the layers without studying the importance of each layer during adaptation. However, the updated LRs, for each layer, do not deviate too much and remain constant throughout.
As the first adaptive LR work for CTTA, LAW [33] hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [21]. To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as,
$s(\theta_{n}^{t};x_{k})=\nabla_{\theta_{n}^{t}}\log(p_{\theta^{t}}(x_{k}))$ (2)
@@
where $x_{k}$ and $\theta_{n}^{t}$ refer to the current test batch and parameters of layer n at time t. Then, the Fisher Information Matrix (FIM) [36], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. As it can be seen, the layer-wise importance (Eq. 2) completely relies on the pseudo-labels to compute the log-likelihood, which is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. Furthermore, the authors propose to capture "domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer during the adaptation, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as,
$\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$ (3)
@@
Based on Eq. 3, this means that LAW [33] puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a "relatively" larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to an imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
@@
Drawing from these primary observations, our method focuses on addressing the following: 1) Based on the prediction uncertainty of the model, we automatically select a subset of layers that would require further adaptation to a task, without relying on the pseudo-labels. 2) As an indicator of the domain shift, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates. We freeze the layers that are not selected.
@@ 4.1 Layer Selection by Quantifying Prediction Uncertainty
CTTA maximizes model performance in the current domain and is vital to assess prediction certainty amidst domain shifts. In the literature on continual learning and TTA, several works utilize FIM [18, 38, 33]. However, the unavailability of true labels, with an unknown distribution, motivates us to capture the prediction uncertainty. Prior literature proposed methods for uncertainty from the classifier output or feature space [12, 14, 24, 25, 20]. Motivated by [15], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence (KL) between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, $\frac{1}{C}$, …., $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C is the total number of classes. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, due to the distribution shift, the KL divergence should be smaller. Let h denote the classifier of the model parameterized by ${\theta^{t}}$. We compute the loss, denoting it as $\mathcal{L}(\theta^{t})$ at time stamp t, as follows,
$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(\hat{h}(x_{k}))||\text{{u}})$ (4)
@@
where $\hat{h}(x_{k})$ is the classifier’s output logits ($h(x_{k})$) smoothed by a temperature coefficient T i.e. $\hat{h}(x_{k})$ = $h(x_{k})$/T. Our intuition is that batches within the same task are of the same distribution and so, smoothing the logits with large values of T will lead to them being indistinguishable. With model parameters $\theta^{t}$, the gradients (${{\nabla\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial{\theta}^{t}}$) can be simplified to,
${{\nabla\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{j=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),j)}{\partial\theta^{t}}$ (5)
@@
On the whole, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq. 5). In test-time adaptation, since ground-truth labels are not present, the gradients are more indicative of familiarity of the model with the batch of data. For batches within the same task, that share the same distribution, the vector norm of Eq. 5 would be higher because of a higher KL divergence since the softmax predictions would be less uniformly distributed. However, if the norm is smaller, this would mean that the softmax outputs are closer to u. Our key idea is to select all the parameters of layers that have the norm of gradients to be less than or equal to a threshold $\eta$. Layers with small gradient norms would require more adaptation for the model to make a confident prediction. Since we freeze the remaining layers, this also helps to alleviate catastrophic forgetting to a large extent by preserving the prior knowledge. And so the scoring function, for each layer, is defined as,
$\mathcal{Z}_{\theta_{n}}=\|{{\nabla\mathcal{L}({\theta}_{n}^{t})}}\|_{1}$ (6)
As mentioned earlier, $\mathcal{Z}_{\theta_{n}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation. We freeze the layers with $\mathcal{Z}_{\theta_{n}}$ $>$ $\eta$ i.e. set their learning rate to 0. This way, in contrast to surgical fine-tuning [21], we automatically select a subset of layers for adaptation to a batch of data.
4.2 Parameter Sensitivity as an Indicator of Domain Shift

### v2

@@ Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Using unlabeled test data, continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains. A highly effective CTTA method involves applying layer-wise adaptive learning rates for selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. This work aims to overcome these limitations by identifying layers for adaptation via quantifying model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity to approximate the domain shift and adjust their learning rates accordingly. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C, demonstrating the superior efficacy of our method compared to prior approaches.
@@ 1 Introduction
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [5, 36]. Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [36]. While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s source knowledge due to adaptations and large parameter updates across tasks.
@@
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In [23, 6, 36, 28], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA [37] includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA [33] introduces the attachment of “meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
@@
More recently, the authors in [15] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically" or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) [24] was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) [27] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
@@
In this paper, we present a novel method called PALM to push the limits of adaptive learning rate methods for CTTA. To address catastrophic forgetting, our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions. Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the Kullback-Leibler (KL) divergence between the model output and a uniform distribution. This approach eliminates dependence on pseudo-labels and enables a more accurate estimation of layer-wise importance. Subsequently, the LRs of the parameters of these selected layers are adjusted based on the degree of domain shift observed in the test data. We use parameter sensitivity as an indicator of domain shift and employ weighted moving averages to aggregate sensitivity from previous test batches. Our approach provides a more accurate estimation of domain shift compared to existing adaptive LR methods, leading to more effective adjustments in LRs.
To summarise, our contributions are as follows: 1) We present PALM, introducing novel improved schemes for CTTA with a specific focus on addressing the key limitations of existing adaptive learning rate methods. 2) Considering the model’s prediction uncertainty at test-time, we automatically select and adapt layers requiring further updates while freezing the rest. We adjust learning rates based on parameter sensitivity to estimate domain shift, making the overall process computationally inexpensive. 3) Our method delivers superior results, via extensive evaluations, in both continual and gradual TTA across three standard benchmark datasets.
@@ 2 Related Works
Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach [12, 34, 36]. TENT [36] optimizes BN parameters by minimizing the Shannon entropy [31] of pseudo labels. BN Stats Adapt (BN-1) [28] replaces the BN statistics with those from corrupted test data, while AdaContrast [3] uses contrastive learning and a memory module to update the entire model. However, these methods address only single domain shifts and overlook continual shifts.
Continual TTA (CTTA): CTTA requires models to adapt to dynamic domain shifts. CoTTA [37] uses a teacher-student setup with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with catastrophic forgetting and error accumulation with time. EcoTTA [33] adds “meta networks" to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting. DSS [38] introduces a dynamic thresholding framework to differentiate between low and high sample quality.
Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work [30, 17, 45, 7]. Depending upon the distribution shift of the target task, in [15], the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive LR methods: Research such as [40, 43] demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers. AutoLR [26] developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training. For CTTA, LAW [24] uses the FIM [27] to assess domain shifts and adjust parameter LRs accordingly. However, using pseudo-labels to calculate FIM may lead to inaccuracies, and accumulating FIMs across domains can result in errors. Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity" of parameters to the adaptation needs of the current batch and domain.
Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity [14, 18, 35, 20]. It suggests that less sensitive parameters require higher LRs for effective adaptation, and vice-versa. Additionally, [16] introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
3 Preliminary
@@ 3.1 Problem Setting
CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time. Each $\mathcal{T}_{i}$ signifies the $i^{th}$ task with $K$ batches, across varying data distributions. At time step t=0, the model is initialized to ${\theta^{s}}$. It is then gradually adapted to each incoming batch $x_{k}$ of $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta^{t}}$. The source dataset is unavailable in this adaptation process due to privacy and storage constraints. With the continuum of tasks, at each $t$, the complete model parameters $\theta^{t}$ consist of a deep tensor with $N$ layers, expressed as $\theta^{t}=\{\theta^{t}_{1},\ldots,\theta^{t}_{N}\}$. Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
3.2 Core Motivations
@@
We provide a detailed view of surgical fine-tuning [15], AutoLR [26], and LAW [24]. It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
@@
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, surgical fine-tuning empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
@@
Inspired by [40, 43], AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$. They hypothesize that initial layers, which extract general features, should have minimal $\delta w_{n}$, while deeper layers extract task-specific features. An adaptive LR optimization scheme constrains weight variations to increase across layers, without assessing each layer’s importance during adaptation. However, the LRs for each layer show minimal deviation and remain consistent throughout the training process.
@@
As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [15]. To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as $s(\theta_{j,n}^{t};x_{k})=\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$, where $p(x_{k})$ is the predicted output. Then, the FIM [27], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. This layer-wise importance scheme completely relies on the pseudo-labels to compute the log-likelihood and is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. The authors also propose to capture “domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
@@
This means that LAW puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively" larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to a possible imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
@@
Drawing from these primary observations, our method focuses on addressing the following: 1) Based on the prediction uncertainty of the model, we automatically select a subset of layers that would require further adaptation to a task, without relying on the pseudo-labels. 2) As a domain shift indicator, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates. We freeze the parameters of the unselected layers.
@@ 4.1 Layer Selection by Quantifying Prediction Uncertainty
CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts. In the literature on continual learning and TTA, several works utilize FIM [13, 29, 24]. However, the unavailability of true labels, with an unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning [15].
@@
Motivated by [11], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where $C$ is the total number of classes of the current task. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, the KL divergence will be small with a large distribution shift. Let $h(\cdot)$ denote the classifier of the model. The loss $\mathcal{L}(\theta^{t})$, to compute the gradients, is computed as follows,
$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(h(x_{k})/T)||\text{{u}})$ (1)
@@
where the classifier’s output logits $h(x_{k})$ is smoothed by a temperature $T$. Let $\hat{h}(x_{k})$ = $h(x_{k})/T$. Our intuition is that since batches within the same task are of the same distribution, smoothing the logits with large values of $T$ will make them indistinguishable. With model parameters $\theta^{t}$, the gradients (${{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial\theta_{j,n}^{t}}$) can be simplified to,
${{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{c=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),c)}{\partial\theta_{j,n}^{t}}$ (2)
@@
Overall, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq. 2). In CTTA, since ground-truth labels are absent, the gradients are more indicative of familiarity of the model with the batch of data. For batches within the same task, that share the same distribution, the vector norm of Eq. 2 would be higher because of a higher KL divergence since the softmax predictions would be less uniformly distributed. However, if the norm is smaller, this would mean that the softmax outputs are closer to u. Our key idea is to select all the parameters of layers with the gradient norms to be less than or equal to a threshold $\eta$. Layers with small gradient norms would require more adaptation for the model to make a confident prediction. We freeze the unselected layers which also helps to alleviate catastrophic forgetting to a large extent by preserving the prior source knowledge. So, the scoring function, for each layer’s parameters, is defined as $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}\|_{1}$. As mentioned earlier, $\mathcal{Z}_{\theta_{j,n}^{t}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation. We freeze the layers with $\mathcal{Z}_{\theta_{j,n}^{t}}$ $>$ $\eta$ i.e., set their LR to 0.
4.2 Parameter Sensitivity as a Domain Shift Indicator